# The cognitive triple-slit experiment

### v1

@@ Abstract
Quantum cognition has made it possible to model human cognitive processes very effectively, revealing numerous parallels between the properties of conceptual entities tested by the human mind and those of microscopic entities tested by measurement apparatuses. The success of quantum cognition has also made it possible to formulate an interpretation of quantum mechanics, called the conceptuality interpretation, which ascribes to quantum entities a conceptual nature similar to that of human concepts. The present work fits into these lines of research by analyzing a cognitive version of single, double, and triple slit experiments. The data clearly show the formation of the typical interference fringes between the slits as well as the embryos of secondary fringes. Our analysis also shows that while quantum entities and human concepts may share a same conceptual nature, the way they manifest it in specific contexts can be quite different. This is also evident from the significant deviation from zero observed for the Sorkin parameter, indicating the presence of strong irreducible third-order interference contributions in human decision.
@@ 1 Introduction
For classical particles, the distribution of impacts when both slits are open can be reconstructed as a uniform average of the distributions with only one slit open. In the quantum case, however, this does not hold due to interference effects. This becomes particularly clear when one observes that the main fringe is located in the center of the two slits, where the probability of an impact is classically the lowest. The formation of an interference fringe corresponds to a region on the screen where the probability is overextended with respect to the classical average (constructive interference), while the regions where no impacts are found are those where the probability is underextended (destructive interference).
@@
Cognitive equivalents of the double-slit experiment also exist [11]. Let us consider an important example. In the 1980s, psychologist James Hampton conducted an experiment in which 40 students were presented with $24$ exemplars of Food and asked to what extent they thought they were related (in terms of membership and typicality) to the concepts: (a) Fruit, (b) Vegetable and (c) Fruit or vegetable [12]. The $24$ exemplars of Food played the same role as the different locations where the trace of an impact can be observed on the detection screen of a double-slit experiment, with the two abstract concepts Fruit and Vegetable playing the role of the two slits. More precisely, choosing a typical exemplar of Fruit is here equivalent to choosing a good example of an impact of an entity passing through the first slit. Similarly, selecting a typical exemplar of Vegetable is equivalent to providing a good example of the impact of an entity passing through the second slit. Finally, selecting a typical exemplar of Fruit or vegetable is equivalent to providing a good example of an impact of an entity passing through both slits, i.e., through the first slit or the second slit.
@@
An expectation based on classical prejudice would tell us that when faced with question (c), about the combination Fruit or vegetable, the probability of selecting a given exemplar of Food would be the uniform average of the probabilities describing the situations of questions (a) and (b), but this is not what Hampton’s data revealed. Similar to the constructive and destructive interference effects in the impact statistics of a typical double-slit experiment, overextensions and underextensions of the probabilities were observed, and when Hampton’s data were represented in a quantum-like way using two wave functions and their superpositions, an interesting interference-like figure was revealed, reminiscent of those obtained in the phenomena of birefringence [11] (see also Figure 4, for a one-dimensional modelization of Hampton’s data).
@@
If we follow this line of reasoning, we can try to provide a cognitive-like interpretation of the double-slit experiment by claiming that the fringes of higher (lower) intensity in the interference pattern correspond to regions that maximize (minimize) doubt about the slits from which the impact came from. In other words, an impact-rich fringe would be a region on the screen that makes it difficult to guess which slit the quantum entity passed through, while it would be exactly the opposite for the regions on the screen with a low density of impacts [11]. Such an interpretation is part of so-called conceptuality interpretation of quantum mechanics, first proposed by Diederik Aerts in 2009 [11] and further developed in his Brussels group [13, 14, 15, 16, 17, 18, 19, 20]. According to it, the behavior of quantum entities appears to be bizarre only if observed on the basis of the prejudice that they are spatiotemporal in nature, for example corpuscles, waves, or even fields. However, if we hypothesize that their nature is conceptual, and that quantum measurements are meaning-driven processes where the measuring apparatuses select outcomes similarly to how cognitive entities respond to questions, then many of the quantum conundrums disappear. Of course, human concepts and the conceptual entities of the microworld should not be confused, in the same way that, for example, electromagnetic and acoustic waves should not be confused, although they share the same undulatory nature.
@@
In this regard, it is important to know that, in the last two decades, a field of research called quantum cognition has emerged, using quantum mechanics, and especially its mathematics, to find new tools for explaining and modeling human cognition, especially how meaning is assigned to concepts and their combinations, and how decision-making processes occur depending on context. It is because of the success of this quantum approach to human modeling that the speculative hypothesis at the basis of the conceptuality interpretation subsequently emerged, according to which this success was not accidental, but due to the fact that quantumness and conceptuality are words that point to a same nature. Indeed, by using the quantum mathematical formalism and the notions that this formalism brings to bear (in particular potentiality and context-dependence), quantum cognition was able to bring to the fore, in the psychological laboratories, all those effects that are typical of quantum entities, such as superposition and interference effects, entanglement, complementarity and indistinguishability [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Positioning itself at the frontier between the conceptuality interpretation and the research in quantum cognition, this article explores to what extent the human mind is able to reproduce a fringe-like pattern (or its embryo) reminiscent of those observed in physics laboratories, when confronted to an experiment that mimics the double and triple-slit experimental situations as closely as possible. More precisely, the article is organized as follows. In Section 2, we recall how the quantum formalism describes the two and three-slit situations. For the latter, we derive the so-called Sorkin parameter, showing that its value is zero when the state of the system is a clear-cut superposition of the different possibilities at play, meaning that there are no genuine third-order interference contributions. In Section 3, we explain why, contrary to what has been initially believed, the quantum formalism also predicts the existence of genuine third-order interferences, albeit in typical experimental settings they are of a much lower magnitude than the second-order ones. In Section 4, we describe the cognitive single, double and triple slit experiments that we have carried out, and in Section 5 we provide a detailed analysis of the collected data. Finally, in Section 6, we offer some concluding reflections.
@@ 2 The quantum triple-slit experiment
where we have assumed that the probabilities of passing through slit-$1$ and slit-$2$ are the same, and $P_{\rm cl}(x|1)$ and $P_{\rm cl}(x|2)$ are the probabilities of observing an impact at point $x$ when only slit-$1$ and slit-$2$ are open, respectively. As mentioned in the previous section, the above average does not hold anymore if the entities are quantum. More precisely, if $\psi(x|1)$ and $\psi(x|2)$ are the wave functions at the detection screen (assumed here to be orthogonal unit vectors), describing the situations where the quantum entity passed through slit-$1$ and slit-$2$ only, respectively, then we know, according to the Born rule, that the associated quantum probabilities are given by the squared modules $P(x|1)=|\psi(x|1)|^{2}$ and $P(x|2)=|\psi(x|2)|^{2}$. When both slits are open, we can apply the prescription saying that when there are alternatives (slit-$1$ or slit-$2$), the probability amplitude is obtained from the normalized sum of the probability amplitudes associated with the alternatives considered separately:
@@
where we have defined the average $\overline{P}(x|123)\equiv\tfrac{1}{3}P(x|1)+\tfrac{1}{3}P(x|2)+\tfrac{1}{3}P(x|3)$, and the two-slit interference contributions $I(x|23)\equiv\Re\,[\psi(x|2)^{*}\psi(x|3)]$, $I(x|13)\equiv\Re\,[\psi(x|1)^{*}\psi(x|3)]$. Inserting (3) in (5), and same for the other two interference terms, after due simplification one gets
@@
called the Sorkin parameter [28], one obtains the condition: $\epsilon(x|123)=0$. So, the Sorkin parameter, an expression consisting only of experimentally accessible probabilities, quantifies possible deviations of the three-slit interference contribution, $I(x|123)\equiv P(x|123)-\overline{P}(x|123)$, from contributions only coming from second-order two-slit interference effects, as is clear that it can also be written
@@
Quantum mechanics predicts that the Sorkin parameter is zero, provided one can interpret the probabilities $P(x|i)$, $i=1,2,3$, as describing the single-slit experiments, i.e., that the prescription (4) is correct, which however is not generally true, as we will discuss in the next section. Note that $\epsilon(x|123)$ is trivially zero in classical mechanics, as all interference contributions are zero.
@@ 3 Third-order interference effects
As first pointed out by Rafael Sorkin [28], the Born rule being quadratic in the wave function it does not allow for genuine third-order interference contributions [29, 30]. From this follows that the Sorkin parameter (7) is zero, which is something that can be tested. The first experiments to measure the Sorkin parameter were performed in 2009 by Urbasi Sinha and colleagues [31, 32, 33]. Their result was that, within their error margins, the nullity of the Sorkin parameter had been verified, and in this sense the prediction of the Born rule confirmed. More precisely, the magnitude of a third-order interference contribution was shown to be less than $10^{-2}$ of the magnitude of the expected second-order interference contributions. Note that any setting that allows for the presence of three mutually exclusive paths can be used to test the nullity of the Sorkin parameter. This was the situation of the experiment performed by Söllner and colleagues [34] with a three-path photonic interferometer, which allowed them to improve by about an order of magnitude the best previous experiments.
@@
Up to this point, it was assumed that Sorkin’s initial analysis and subsequent experiments were to be interpreted as confirming the peculiarity of the Born rule in not predicting genuine third-order interference contributions. But things were not as they appeared. As pointed out in [35], it was incorrect to assume that it was always possible to write the wave function on the detection screen, when all three slits are open, as the superposition (4) of contributions due only to the three alternatives in play, i.e., to passing through only one of the three slits. One way to understand this is to observe that, due to the phenomenon of the spreading of the wave-packet, it is not possible to decompose a wave-function on the screen as a superposition of wave-functions uniquely associated with the single-slit situations. In fact, during its evolution in time, the support of any wave function will inevitably always extend to touch with its tails all the slits. This can be expressed in a more specific way by considering a Feynman path-integral approach [36], which involves integration over all possible paths, including those passing through several slits (forming loops for example) before reaching the detection screen [37], and this is another way of saying that the wave function is not uniquely decomposable as a superposition of contributions attributable to individual slits. While these looped paths contribute less to the overall probability than paths closer to classical trajectories, they cannot be neglected. In other words, contrary to the initial belief, a non-zero value of the Sorkin parameter was not to be considered as a violation of the Born rule [35, 38, 39, 40]. To explain it yet another way, Yabuki emphasized that when we close slits we modify the Hamiltonian. Thus, in a double-slit problem, the solutions with only one slit open are not directly related to the solution with both slits open, since they correspond to different physical contexts, so the superposition principle cannot be applied naively [37]. And of course the same is true for a three-slit experiment.
@@
It should be emphasized that even in the two-slit case there are exotic paths that pass through both slits and are contained in the second-order interference contributions. This means that a non-zero Sorkin parameter accounts for more complex paths that interact with all three slits. If they usually constitute a minimal correction [39], they become more relevant as the operating wavelength increases with respect to the slit spacing. This is because the longer the wavelength, the greater the overlap between the wave functions of the individual slits, which gives more weight to the non-classical paths interacting with all the slits. This explains why the 2018 experiment by Govindaraj Rengaraj et al., that measured a non-zero Sorkin parameter, worked in the microwave regime [43]. So, in quantum mechanics, when we are in frequency regimes (or experimental designs) where the exotic paths contributions are not relevant, there are no genuine third-order interference effects, but when we move from the physical to the psychological laboratory, things may change, and it is one of the aims of this article to investigate this possibility.
@@
We are not aware of any tests for possible violations of the nullity of the Sorkin parameter in experiments with combinations of three concepts, and it is certainly our intention, in the near future, to carry out such an experiment. However, we can certainly reason as follows. The three two-word disjunctions Fruit or vegetable, Fruit or flavoring, and Vegetable or flavoring express different situations of doubt about the classification of a given exemplar of Food. Imagine that an exemplar cannot be conveniently classified as either Fruit, Vegetable, or Flavoring. This will lead the corresponding two-word conjunctions receiving higher scores, i.e., being overextended, precisely because they express a situation of doubt. However, the three-word conjunction Fruit or vegetable or flavoring also conveys a meaning of doubt about how to classify the exemplar in question, and undoubtedly does so differently from the two-word conjunctions, since it explicitly contains all three words. The question that naturally arises is whether this three-word (three-slit-like) expression of doubt allows, in real experiments, meanings to be derived that are in no way deducible from two-word (two-slit-like) expressions of doubt. If this is the case, the equality $\epsilon(x|123)=0$ is going to be violated. But will it be a weak violation, as is usually the case in quantum mechanics, or a strong one? To answer, we have designed and conducted a cognitive experiment in which we simulated single, double and triple-slit experiments, asking participants to directly place impact points on the equivalent of a detection screen. Of course, this required simplifying and converting the experiments into appropriate cognitive terms, as we will explain in the next section.

### v2

@@ Abstract
Quantum cognition has made it possible to model human cognitive processes very effectively, revealing numerous parallels between the properties of conceptual entities tested by the human mind and those of microscopic entities tested by measurement apparatuses. The success of quantum cognition has also made it possible to formulate an interpretation of quantum mechanics, called the conceptuality interpretation, which ascribes to quantum entities a conceptual nature similar to that of human concepts. The present work fits into these lines of research by analyzing a cognitive version of single-slit, double-slit, and triple-slit experiments. The data clearly show the formation of the typical interference fringes between the slits, as well as the embryos of secondary fringes. Our analysis also shows that while quantum entities and human concepts may share a same conceptual nature, the way they manifest it in specific contexts can be quite different. This is also evident from the significant deviation from zero observed for the Sorkin parameter, indicating the presence of strong irreducible third-order interference contributions in human decision.
@@ 1 Introduction
For classical particles, the distribution of impacts when both slits are open can be reconstructed as a uniform average of the distributions with only one slit open at a time. In the quantum case, however, this does not hold, due to interference effects. This becomes particularly clear when one observes that the main fringe is located in the center of the two slits, where the probability of an impact is classically the lowest; see Fig. 1. The formation of an interference fringe corresponds to a region on the screen where the probability is overextended with respect to the classical average (constructive interference), while the regions where no impacts are found are those where the probability is underextended (destructive interference).
Figure 1: A schematic representation of the two-slit experiment in the case of classical and quantum entities. In the classical case (top right), the pattern on the screen can be deduced from a uniform average of the two patterns when the two slits are opened one at a time. In the quantum case (bottom right), the fringe pattern cannot be deduced from those obtained when only one slit remains open at a time.
@@
Cognitive equivalents of the double-slit experiment also exist [11]. Let us consider an important example. In the 1980s, psychologist James Hampton conducted an experiment in which 40 students were presented with $24$ exemplars of Food and asked to what extent they thought they were related (in terms of membership and typicality) to the concepts: (a) Fruit, (b) Vegetable, and (c) Fruit or vegetable [12]. The $24$ exemplars of Food played the same role as the different locations where the trace of an impact can be observed on the detection screen of a double-slit experiment, with the two abstract concepts Fruit and Vegetable playing the role of the two slits. More precisely, choosing a typical exemplar of Fruit is here equivalent to choosing a good example of an impact of an entity passing through the first slit. Similarly, selecting a typical exemplar of Vegetable is equivalent to providing a good example of the impact of an entity passing through the second slit. Finally, selecting a typical exemplar of Fruit or vegetable is equivalent to providing a good example of an impact of an entity passing through both slits, i.e., through the first slit or the second slit.
@@
An expectation based on classical prejudice would tell us that when faced with question (c), about the combination Fruit or vegetable, the probability of selecting a given exemplar of Food would be the uniform average of the probabilities describing the situations of questions (a) and (b), but this is not what Hampton’s data revealed. Similar to the constructive and destructive interference effects in the impact statistics of typical double-slit experiments, overextensions and underextensions of the probabilities were observed, and when Hampton’s data were represented in a quantum-like way, using two Gaussian wave functions and their superpositions, an interesting interference-like figure was revealed, reminiscent of those obtained in the phenomena of birefringence [11]; see Figure 2 see also Figure 10, for a one-dimensional modelization of Hampton’s data).
Figure 2: (color online) The first two figures on the left represent, respectively, the experimental probabilities of selecting a typical exemplar of Fruit, with Apple being the most frequently chosen, and the experimental probabilities of selecting a typical exemplar of Vegetable, with Broccoli being the most frequently chosen. The figure at the top right represents the probabilities obtained by considering a uniform average of the probabilities for Fruit and for Vegetable. This average, typical for classical entities, differs greatly from the data observed by Hampton, when the conceptual combination Fruit or vegetable was considered. The figure at the bottom right represents the experimental probabilities of selecting a typical exemplar of Fruit or vegetable, modeled using the quantum superposition principle. It exhibits interference phenomena similar to those of birefringence. For more information on how these curves were constructed, see Section 6 and [11].
@@
If we follow this line of reasoning, we can try to provide a cognitive-like interpretation of the double-slit experiment by claiming that the fringes of higher (lower) intensity in the interference pattern correspond to regions that maximize (minimize) doubt about the slits from which the impact came from. In other words, an impact-rich fringe would be a region on the screen that makes it difficult to guess which slit the quantum entity passed through, while it would be exactly the opposite for the regions on the screen with a low density of impacts [11]. Such an interpretation is part of the so-called conceptuality interpretation of quantum mechanics, first proposed by Diederik Aerts in 2009 [11] and further developed in his Brussels group [13, 14, 15, 16, 17, 18, 19, 20]. According to it, the behavior of quantum entities appears to be bizarre only if observed on the basis of the prejudice that they are spatiotemporal in nature, for example corpuscles, waves, or even fields. However, if we hypothesize that their nature is conceptual, and that quantum measurements are meaning-driven processes where the measuring apparatuses select outcomes similarly to how cognitive entities respond to questions, then many of the quantum conundrums disappear. Of course, human concepts and the conceptual entities of the microworld should not be confused, in the same way that, for example, electromagnetic and acoustic waves should not be confused, although they share the same undulatory nature.
@@
In this regard, it is important to know that, in the last two decades, a field of research called quantum cognition has emerged, using quantum mechanics, and especially its mathematics, to find new tools for explaining and modeling human cognition, especially how meaning is assigned to concepts and their combinations, and how decision-making processes occur depending on context. It is because of the success of this quantum approach to human modeling that the speculative hypothesis at the basis of the conceptuality interpretation subsequently emerged, according to which this success was not accidental, but due to the fact that quantumness and conceptuality are notions that point to a same nature. Indeed, by using the quantum mathematical formalism and the notions that this formalism brings to bear (in particular, potentiality and context-dependence), quantum cognition was able to bring to the fore, in the psychological laboratories, all those effects that are typical of quantum entities, such as superposition and interference effects, entanglement, complementarity and indistinguishability [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Positioning itself at the frontier between the conceptuality interpretation and the research in quantum cognition, this article explores to what extent the human mind is able to reproduce a fringe-like pattern (or its embryo) reminiscent of those observed in physics laboratories, when confronted to an experiment that mimics the double and triple-slit experimental situations as closely as possible. More precisely, the article is organized as follows. In Section 2, we recall how the quantum formalism describes the two and three-slit situations. For the latter, we derive the so-called Sorkin parameter, showing that its value is zero when the state of the system is a clear-cut superposition of the different possibilities at play, meaning that there are no genuine third-order interference contributions. In Section 3, we explain why, contrary to what has been initially believed, the quantum formalism also predicts the existence of genuine third-order interferences, albeit in typical experimental settings they are of a much lower magnitude than the second-order ones. In Section 4, we describe the cognitive single-slit, double-slit and triple-slit experiments that we have carried out, and in Section 5 we provide a detailed analysis of the collected data. Finally, in Section 6, we offer some concluding reflections.
@@ 2 The quantum triple-slit experiment
where we have assumed that the probabilities of passing through slit-$1$ and slit-$2$ are the same, and $P_{\rm cl}(x|1)$ and $P_{\rm cl}(x|2)$ are the probabilities of observing an impact at point $x$ when only slit-$1$ and slit-$2$ are open, respectively. As mentioned in the previous section, the above average does not hold anymore if the entities are quantum. More precisely, if $\psi(x|1)$ and $\psi(x|2)$ are the wave functions at the detection screen (assumed here to be orthogonal unit vectors), describing the situations where the quantum entity passed through slit-$1$ and slit-$2$ only, respectively, then we know, according to the Born rule, that the associated quantum probabilities are given by the squared modules $P(x|1)=|\psi(x|1)|^{2}$ and $P(x|2)=|\psi(x|2)|^{2}$. When both slits are open, we can apply the prescription saying that when there are alternatives (slit-$1$ or slit-$2$), the probability amplitude, $\psi(x|12)$, is obtained from the normalized sum of the probability amplitudes associated with the alternatives considered separately:
@@
where we have defined the (classical) average $\overline{P}(x|123)\equiv\tfrac{1}{3}P(x|1)+\tfrac{1}{3}P(x|2)+\tfrac{1}{3}P(x|3)$, and the two-slit interference contributions $I(x|23)\equiv\Re\,[\psi(x|2)^{*}\psi(x|3)]$, $I(x|13)\equiv\Re\,[\psi(x|1)^{*}\psi(x|3)]$. Inserting (3) in (5), and same for the other two interference terms, after due simplification one gets:
@@
called the Sorkin parameter [28], one obtains the condition: $\epsilon(x|123)=0$, for all $x$. So, the Sorkin parameter, an expression consisting only of experimentally accessible probabilities, quantifies possible deviations of the three-slit interference contribution, $I(x|123)\equiv P(x|123)-\overline{P}(x|123)$, from contributions only coming from second-order two-slit interference effects, as is clear that it can also be written
@@
Quantum mechanics predicts that the Sorkin parameter is zero, provided one can interpret the probabilities $P(x|i)$, $i=1,2,3$, as describing the single-slit experiments, i.e., that the prescription (4) is correct, which however is not generally true, as we will discuss in the next section. Note that $\epsilon(x|123)$ is trivially zero in classical mechanics, as all interference contributions are then zero.
@@ 3 Third-order interference effects
As first pointed out by Rafael Sorkin [28], the Born rule being quadratic in the wave function it does not allow for genuine third-order interference contributions [29, 30]. From this follows that the Sorkin parameter (7) is zero, which is something that can be experimentally tested. The first experiments to measure the Sorkin parameter were performed in 2009 by Urbasi Sinha and colleagues [31, 32, 33]. Their result was that, within their error margins, the nullity of the Sorkin parameter had been verified, and in this sense the prediction of the Born rule confirmed. More precisely, the magnitude of a third-order interference contribution was shown to be less than $10^{-2}$ of the magnitude of the expected second-order interference contributions. Note that any setting that allows for the presence of three mutually exclusive paths can be used to test the nullity of the Sorkin parameter. This was the situation of the experiment performed by Söllner and colleagues [34] with a three-path photonic interferometer, which allowed them to improve by about an order of magnitude the best previous experiments.
@@
Up to this point, it was believed that Sorkin’s initial analysis and subsequent experiments were to be interpreted as confirming the peculiarity of the Born rule in not predicting genuine third-order interference contributions. But things were not as they appeared. As pointed out in [35], it was incorrect to assume that it is always possible to write the wave function on the detection screen, when all three slits are open, as the superposition (4) of contributions due only to the three alternatives in play, i.e., to passing through only one of the three slits. One way to understand this is to observe that, due to the phenomenon of the spreading of the wave-packet, it is not possible to decompose a wave-function on the screen as a superposition of wave-functions uniquely associated with the single-slit situations. In fact, during its evolution in time, the support of any wave function will inevitably always extend to touch with its tails all the slits. This can be expressed in a more specific way by considering a Feynman path-integral approach [36], which involves integration over all possible paths, including those passing through several slits (forming loops for example) before reaching the detection screen [37], and this is another way of saying that the wave function is not uniquely decomposable as a superposition of contributions attributable to individual slits. While these looped paths contribute less to the overall probability than paths closer to classical trajectories, they cannot be neglected. In other words, contrary to the initial belief, a non-zero value of the Sorkin parameter was not to be considered as a violation of the Born rule [35, 38, 39, 40]. To explain it yet another way, Yabuki emphasized that when we close slits we modify the Hamiltonian. Thus, in a double-slit problem the solutions with only one slit open are not directly related to the solutions with both slits open, since they correspond to different physical contexts, so the superposition principle cannot be applied naively [37]. And of course the same is true for a three-slit experiment.
@@
Figure 3: Three examples of exotic paths that cross all three slits. See the analysis in [37].
It should be emphasized that even in the two-slit case there are exotic paths that pass through both slits and are contained in the second-order interference contributions. This means that a non-zero Sorkin parameter accounts for complex paths that interact with all three slits. If they usually constitute a minimal correction [39], they become more relevant as the operating wavelength increases with respect to the slit spacing. This is because the longer the wavelength, the greater the overlap between the wave functions of the individual slits, which gives more weight to the non-classical paths interacting with all the slits. This explains why the 2018 experiment by Govindaraj Rengaraj et al., that measured a non-zero Sorkin parameter, worked in the microwave regime [43]. So, in quantum mechanics, when we are in frequency regimes (or experimental designs) where the exotic paths contributions are not relevant, there are no genuine third-order interference effects, but when we move from the physical to the psychological laboratory, things may change, and it is one of the aims of this article to investigate this possibility.
@@
We are not aware of any tests for possible violations of the nullity of the Sorkin parameter in experiments with combinations of three concepts, and it is certainly our intention, in the near future, to carry out such an experiment. However, we can certainly reason as follows. The three two-word disjunctions, Fruit or vegetable, Fruit or flavoring and Vegetable or flavoring, express different situations of doubt about the classification of a given exemplar of Food. Imagine that an exemplar cannot be conveniently classified as either Fruit, Vegetable, or Flavoring. This will lead the corresponding two-word conjunctions receiving higher scores, i.e., being overextended, precisely because they express a situation of doubt. However, the three-word conjunction, Fruit or vegetable or flavoring, also conveys a meaning of doubt about how to classify the exemplar in question, and undoubtedly does so differently from the two-word conjunctions, since it explicitly contains all three words. The question that naturally arises is whether this three-word (three-slit-like) expression of doubt allows, in real experiments, meanings to be derived that are in no way deducible from two-word (two-slit-like) expressions of doubt. If this is the case, the equality $\epsilon(x|123)=0$ is going to be violated. But will it be a weak violation, as is usually the case in quantum mechanics, or a strong one? The cognitive test we have conducted, which is described in the next section and simulates single-slit, double-slit and triple-slit experiments, asking participants to place impact points directly on the equivalent of a detection screen, will allow us to offer an initial response to this question.